OpenStudy (anonymous):

a particle moves with constant acceleration for 3 seconds, during which time it travels 27m. it then moves with constant velocity and during the nest 5 seconds travels 60m. so, whats the acceleration of the particle in the first 3 seconds? and what was the initial velocity of it?? thanks :)

OpenStudy (anonymous):

6 m/s

OpenStudy (anonymous):

Can you see how?

OpenStudy (anonymous):

Work out from the second part s=vt 60=v*5 v=12m/s Now for the first part s=(v+u)*t/2 27=(12+u)*3/2 u=6

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

My pleasure as always........

OpenStudy (anonymous):

can you also help me with two other questions please??? cause you seems to understand this

OpenStudy (anonymous):

Where are the questions?